OpenStudy (anonymous):

(Find one counter example to show that each conjecture is false.) The product of two positive numbers is greater than either number.

OpenStudy (dangerousjesse):

Any number and zero would make this false.

OpenStudy (anonymous):

so would any two numbers less than one

OpenStudy (anonymous):

okay I think I understand

OpenStudy (nikato):

Can't work because 0 is not positive

OpenStudy (dangerousjesse):

Yep.. There's an infinite amount of options.

OpenStudy (anonymous):

so like 2+4 >0 ??

OpenStudy (dangerousjesse):

Ah, very good point haha.

OpenStudy (dangerousjesse):

Product=Multiplication

OpenStudy (anonymous):

@KellySund product means multiply

OpenStudy (anonymous):

oh okay!

OpenStudy (dangerousjesse):

Are \(8>2\) and \(8>4\) both true?

OpenStudy (anonymous):

yes

OpenStudy (dangerousjesse):

Then your example was true :)

OpenStudy (anonymous):

Okay~!

OpenStudy (aum):

\(\large \frac 12 * \frac 14 = \frac 18\\ \large \frac 18 < \frac 12 \\ \large \frac 18 < \frac 14 \)